OpenStudy (anonymous):

A=lw 3x^3= length and 2x^4= width I got 86 is this correct find the area of a rectangle

OpenStudy (phi):

It depends on what x is

OpenStudy (texaschic101):

6x^7 ??

OpenStudy (phi):

yes 6x^7 is the area (length times width) to get a number like 86, we need to put in a number for x

OpenStudy (anonymous):

Here is what it says: " Using the formula A=lw, find the area of the rectangle. 3x^3 and 2x^4. Find the area? so Imultiplied the length and width including the exponents and this is how I got 86.

OpenStudy (texaschic101):

I still can't figure out how your getting 86 ???

OpenStudy (anonymous):

6x^7 = 279936

OpenStudy (phi):

when you multiply 3x^3 times 2x^4 you cannot get a nice number. you can multiply 3*2 to get 6 x^3 means x*x*x x^4 means x*x*x*x multiply them, you get: x*x*x*x*x*x*x which is abbreviated x^7 the answer is 6*x^7 or just 6x^7

OpenStudy (texaschic101):

3x * 2x = 6x x^3 * x^4 = x^7 put them together and you get 6x^7

OpenStudy (texaschic101):

oops...3 * 2 = 6

OpenStudy (phi):

no 3^3 means 3*3*3 3x^3 means 3*x*x*x the answer will have x's

OpenStudy (phi):

3* x^3 * 2* x^4 you can change the order of the multiplies to 3*2* x^3 * x^4 you can simplify 3*2 to just 6 6 * x^3 * x^4 and as shown above, you can simplify x^3 * x^4 to x^7 6 * x^7 normally, you don't bother to show the multiply (the * here) and write this as 6x^7 which is short for 6*x*x*x*x*x*x*x

OpenStudy (anonymous):

thank you so much!